OpenStudy (mathhater):

Write the formula given the following arithmetic sequences. 4, 2, 0, -2

OpenStudy (anonymous):

a = 2 d = 2 - 4 = -2 So, for last term, tn = a + (n - 1)*d = 2 + (n - 1)*(-2) tn = 2 + 2n - 2 = 2n.. So, the formula given to this Arithmetic Progression is : tn = 2n...